# Woosh: A Sound Effects Foundation Model

### 13 Apr 2026 v2

@@ 2.2 Data
  • Freesound [15] - A dataset consisting of Creative Commons licensed sound clips collected on the freesound.org online platform. Only audio with public domain (CC0), CC-BY, and CC-Sampling+ was used. The dataset contains a total of 370 k audio files, with the vast majority sampled at least at 44.1 kHz. We kept the original sample descriptions from Freesound.
@@ 3.2 Data
To train the Woosh-CLAP-Public model, we used the Freesound, AudioCaps, and WavCaps datasets described in Section 2.2. The Woosh-CLAP-Private model was trained using only the same mix of commercial studio-quality sound effect libraries as described in the same section. Compared to the public datasets, the captions provided by these internal datasets were clearly less verbose, keyword-like, and using specific technical vocabulary and abbreviations. We used a large language model (LLM) to generate alternative descriptions that were more verbose and closer to natural language.
@@ 3.3 Training
where $N$ is the batch size and $\tau$ is a temperature value. The numerators are computed for the positive pairs, i.e., the diagonal elements in the cosine similarity matrix of Figure 3, and the denominators are computed for all pairs, i.e., all cosine similarity elements.
@@ 4.2 Data
To train the Woosh-Flow-Public model we use the Freesound, AudioCaps, and WavCaps datasets described in Section 2.2. To train the Woosh-Flow-Private model we use the same mix of commercial studio-quality sound effect libraries described in the same section. Pre-processing consisted of resampling the audio to 48 kHz. Audio samples were pre-segmented for non-silence using an energy-based detector, joining contiguous segments closer than 2 seconds. Random 5-second long chunks from a random segment were taken for training and testing.
@@ 5.4 Distillation
To accelerate inference for the video-to-audio model, we apply a distillation process highly analogous to the text-to-audio MeanFlow (MF) approach detailed in Section 4.4. We refer to this efficient model as Woosh-DVFlow. By leveraging the teacher model’s predictions alongside CFG distillation, we successfully reduce the required NFEs at inference time from 100 to 4 while maintaining a high-fidelity audio generation.
@@ 5.5 Evaluation
Table 4 presents the quantitative evaluation of our models compared to the MMAudio-M baseline. On the high-quality FoleyBench dataset, our Woosh-VFlow architecture demonstrates strong audio generation capabilities despite having fewer parameters than the baseline (413 M vs. 621 M). Woosh-VFlow consistently achieves better audio quality, yielding lower FD and KL divergence scores in both text-conditioned and video-only settings. Furthermore, our distilled model, Woosh-DVFlow, successfully maintains highly competitive metrics on FoleyBench while reducing the required inference steps from 100 to 4. We also observe a substantial performance gap in favor of Woosh-VFlow on the OGameData dataset. However, it should be noted that our model was explicitly trained on in-domain gameplay videos (OGameData250k), which is expected to provide a significant advantage in this specific domain compared to the MM-Audio.

### 29 Apr 2026 v3

@@ 2.2 Data
  • Freesound [15] - A dataset consisting of Creative Commons licensed sound clips collected on the freesound.org online platform. Only audio with public domain (CC0), $C$-BY, and $C$-Sampling+ was used. The dataset contains a total of 370 k audio files, with the vast majority sampled at least at 44.1 kHz. We kept the original sample descriptions from Freesound.
@@ 3.2 Data
To train the Woosh-CLAP-Public model, we used the Freesound, AudioCaps, and WavCaps datasets described in Section $2.$ The Woosh-CLAP-Private model was trained using only the same mix of commercial studio-quality sound effect libraries as described in the same section. Compared to the public datasets, the captions provided by these internal datasets were clearly less verbose, keyword-like, and using specific technical vocabulary and abbreviations. We used a large language model (LLM) to generate alternative descriptions that were more verbose and closer to natural language.
@@ 3.3 Training
where NN is the batch size and $\tau$ is a temperature value. The numerators are computed for the positive pairs, i.e., the diagonal elements in the cosine similarity matrix of Figure 3, and the denominators are computed for all pairs, i.e., all cosine similarity elements.
@@ 4.2 Data
To train the Woosh-Flow-Public model we use the Freesound, AudioCaps, and WavCaps datasets described in Section $2.$ To train the Woosh-Flow-Private model we use the same mix of commercial studio-quality sound effect libraries described in the same section. Pre-processing consisted of resampling the audio to 48 kHz. Audio samples were pre-segmented for non-silence using an energy-based detector, joining contiguous segments closer than 2 seconds. Random 5-second long chunks from a random segment were taken for training and testing.
@@ 5.4 Distillation
To accelerate inference for the video-to-audio model, we apply a distillation process highly analogous to the text-to-audio MeanFlow (MF) approach detailed in Section $4.$ We refer to this efficient model as Woosh-DVFlow. By leveraging the teacher model’s predictions alongside CFG distillation, we successfully reduce the required NFEs at inference time from 100 to 4 while maintaining a high-fidelity audio generation.
@@ 5.5 Evaluation
Table 4 presents the quantitative evaluation of our models compared to the MMAudio-M baseline. On the high-quality FoleyBench dataset, our Woosh-VFlow architecture demonstrates strong audio generation capabilities despite having fewer parameters than the baseline (413 M vs. 621 M). Woosh-VFlow consistently achieves better audio quality, yielding lower FD and KL divergence scores in both text-conditioned and video-only settings. Furthermore, our distilled model, Woosh-DVFlow, successfully maintains highly competitive metrics on FoleyBench while reducing the required inference steps from 100 to 4. We also observe a substantial performance gap in favor of Woosh-VFlow on the OGameData dataset. However, it should be noted that our model was explicitly trained on in-domain gameplay videos (OGameData250k), which is expected to provide a significant advantage in this specific domain compared to the $M$-Audio.